# On the Effectiveness of Methods and Metrics for Explainable AI in Remote Sensing Image Scene Classification

### v1

@@ I Introduction
The development of accurate methods for remote sensing (RS) image scene classification is one of the most important research topics in RS. In recent years, deep learning (DL)has emerged as the dominant approach for scene classification problems due to the high capability of DL models to extract and exploit the complex spatial and spectral content of RS images. RS image scenes can contain multiple land-use land-cover (LULC) classes and thus can be simultaneously associated with different class labels (i.e. multi-labels). Accordingly, the existing models are in general capable of assigning multiple LULC class labels to each RS image scene in an archive [1, 2]. Despite their success, DL models often remain enigmatic black boxes, with almost no human understanding of their internal processes [3]. This lack of transparency raises concerns as decision-makers increasingly rely on these black box systems [4]. New regulations for critical applications, such as the General Data Protection Regulation [5] and the Artificial Intelligence Act [6], require interpretability to ensure transparent decisions. Beyond regulatory concerns, effective interpretability can also improve model performance by addressing issues such as limited or biased training data, outliers, vulnerability to adversarial attacks, and overfitting [4].
@@
To this end, explainable artificial intelligence (xAI)has emerged as a research area focused on revealing the factors behind DL model predictions. Although xAI methods are widely used in the CV community [7], they have only recently gained attention in RS. For a comprehensive overview of xAI in RS we refer the reader to [8]. The most prominent type of explanation methods for RS scene classification problems are feature attribution methods [8]. These methods assign importance scores to input features based on their contribution to the prediction of a model[9, 10, 11, 12, 13]. While some feature attribution methods have been adapted to RS scene classification[14, 15], most studies use feature attribution methods that were originally designed for image classification tasks on natural images [16, 17, 18].
@@
This raises the question of whether they are applicable to RS image scenes, as they exhibit different characteristics compared to natural images and can even be considered a distinct modality [19]. To answer this question, it is essential to assess the quality of the generated explanations. Existing evaluation methods can be categorized into two groups: 1) qualitative assessments, which include domain expert evaluations or user studies; and 2) quantitative assessments, which employ explanation metrics [20]. Recently, Höhl et al. [8] reviewed approximately 200 RS publications, finding that the vast majority evaluated qualitative, while only 16 studies included quantitative evaluation using explanation metrics. This reliance on qualitative assessment raises concerns, as studies from the CV community have shown that qualitative assessment is insufficient for robust verification of explanation methods [21, 22].
@@
One of the key reasons for the limited use of quantitative evaluations in RS is the absence of clear guidelines for assessing explanation quality in this domain. The need for such guidelines is underscored by the inherent difficulty of quantitative evaluation since explanation metrics depend heavily on their parametrization [23, 22]. Moreover, the lack of ground truth for the explanations renders them unverifiable [24], requiring the use of a wide range of metrics for a more comprehensive evaluation [25]. In many RS studies only a small part of this range is analyzed, often focusing on single aspects, such as localization capability [17], while only a few studies have evaluated a broader variety of metrics [26, 27]. In particular, none of these studies have addressed the fact that most explanation metrics have been originally designed for natural images and can be limited when applied to RS scene classification.
@@
To address these issues, we present the first unified evaluation of the explanation methods and metrics for RS image scene classification. We assess five widely used feature attribution methods: Occlusion [9], LIME [10], LRP [11], DeepLIFT [12] and GradCAM [13]; and evaluate ten explanation metrics spanning five categories: faithfulness (FE [28], IROF [29]), robustness (AS [30], LLE [28]), localization (RRA [31], TKI [32]), complexity (SP [33], CO [34]) and randomization-based (MPRT [35], RL [36]). These methods and metrics are selected for their popularity in RS and to cover a broad spectrum of evaluation criteria. We conduct a methodological analysis to examine whether existing explanation methods and metrics are suitable for RS image scene classification. Then, we perform an extensive experimental evaluation using MetaQuantus [24] to assess the reliability and effectiveness of explanation metrics under minor perturbations and disruptive changes, supporting our methodological findings with experimentally validation.
@@
  • Based on our methodological and experimental findings, we develop practical guidelines to assist RS practitioners in selecting appropriate explanation methods and evaluation metrics for RS image scene classification and recommend suitable explanation methods and metrics. Our pipeline can be reproduced and adapted for different datasets or classification tasks in RS, providing a valuable resource for the community.
The remainder of this paper is organized as follows. Section II and Section III provide a methodological evaluation of the effectiveness of explanation methods and explanation metrics for RS image scene classification, respectively. In Section IV we describe the datasets and experimental setup. The results of our experiments are presented in Section V. Finally, Section VI draws the conclusion of the work.
Figure 1: The evaluation protocol for selecting appropriate explanation methods and metrics for RS image scene classification. It involves generating explanations using various feature attribution methods, estimating their quality using selected metrics, assessing the reliability of the metric with MetaQuantus, and using the most reliable metrics to select the most suitable explanation methods.
@@ II Explanation methods for RS image scene classification
In this section, we methodologically analyze the explanation methods to address the unique challenges posed by the RS images. Most xAI research has focused on natural images, which differ significantly from RS images [19]. To this end, we first analyze a range of explanation methods commonly applied in RS scene classification regarding these differences. We focus on local post-hoc explanation methods that generate feature attribution maps, as they are most commonly used in RS [8]. Feature attribution methods can be further grouped into backpropagation-based and perturbation-based methods [37]. Perturbation-based methods alter model inputs to observe changes in the model output, while backpropagation methods propagate values from the model output to the input space, typically gradients or relevancy scores. We consider three categories: i) perturbation-based methods, ii) relevancy-propagation-based methods and iii) gradient-propagation-based methods. To provide a comprehensive overview, our analysis includes five popular feature attribution methods covering all categories: 1) two perturbation-based methods (Occlusion [9], LIME [10]); 2) two relevancy-propagation-based methods (LRP [11], DeepLIFT [12]) and 3) one gradient-propagation-based method (GradCAM [13]). For each category, we first provide a general description, followed by a definition of the explanation methods and a discussion of the limitations considering the differences between RS and natural images.
Let us consider a RS image scene classification problem where a DL model $f$ is trained on a RS image archive $\bm{X}_{\mathrm{tr}}$ made up of $N$ pairs: $(\bm{x}_{i},y_{i})_{i=1}^{N}$ for single‐label classification, or $(\bm{x}_{i},\bm{y}_{i})_{i=1}^{N}$ for multi‐label classification. For simplicity, we omit the single-label notation. Each image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ represents a scene captured by an airborne or satellite sensor, where $C$ denotes the spectral bands and $H,W$ correspond to spatial dimensions. The model $f$ maps an input image $\bm{x}$ to a predicted set of labels:
$f(\bm{x})=\hat{\bm{y}}$ (1)
where $\hat{\bm{y}}$ represents the predicted class probabilities or binary assignments. To interpret the reasoning behind the prediction of a model $f$ for an input $\bm{x}$, post-hoc explanation methods, such as feature attribution methods $\Phi$, can be applied:
@@
II A Perturbation-based methods
II A 1 Background
Perturbation-based explanation methods explain model predictions by measuring changes in the output between the original and perturbed inputs. Let $\mathbb{X}\subseteq\mathbb{R}^{H\times W\times C}$ be the space of valid inputs (e.g., images) and let $\Theta$ be a set of perturbation parameters (e.g., masks, noise levels, patch locations). A perturbation function $\mathcal{P}_{\mathbb{X}}$ is defined as:
$\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\Theta\;\to\;\mathbb{X},$ (3)
such that for any $\bm{x}\in\mathbb{X}$ and any $\theta\in\Theta$, the output
$\tilde{\bm{x}}\;=\;\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ (4)
remains in $\mathbb{X}$. In other words, $\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ modifies $\bm{x}$ according to the perturbation parameter $\theta$, while still preserving the overall structure such that $\tilde{\bm{x}}$ is a valid point in the same space. We define the category of perturbation-based methods $\Phi_{\text{Perturbation}}$ as:
$\Phi_{\text{Perturbation}}(f,\bm{x},\mathcal{P}_{\mathbb{X}},\theta)=|f(\bm{x})-f(\mathcal{P}_{\mathbb{X}}(\bm{x},\theta))|.$ (5)
Common perturbation functions are blurring [38], introducing noise [30], or replacing parts of the image [39]. For perturbation-based explanation methods, we consider Occlusion [9] and LIME [10].
The Occlusion method [9] is defined as follows. Let each $\theta\in\Theta$ specify a region (mask) $\Omega\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ and a constant perturbation baseline $\alpha\in\mathbb{R}$. We define the perturbation function:
@@ II A 1 Background
The LIME $\Phi_{\mathrm{LIME}}$ method explains a black-box model $f$ by locally approximating it with a simple, interpretable model trained on binary-encoded perturbations of the original input. Let $S=\{S_{1},S_{2},\dots,S_{M}\}$ be a segmentation of $\bm{x}$ into $M$ superpixels. Denote by $\bm{z}\in\{0,1\}^{M}$ a binary vector that indicates which superpixels are present (1) or not present (0). We define the perturbation function:
@@
and $\alpha\in\mathbb{R}^{C}$ is a constant perturbation baseline. By toggling each superpixel via $\bm{z}_{m}\in\{0,1\}$, we create perturbed samples in the neighborhood of $\bm{x}$. LIME estimates a simple, interpretable model $g$ (e.g., a sparse linear model) that approximates $f$ on these perturbations near $\bm{x}$. Formally, let:
@@
$\Phi_{\mathrm{LIME}}(\bm{x})\;=\;\operatorname{argmin}_{g\in\mathcal{G}}\Bigl{[}\mathcal{L}\bigl{(}f,g,\Pi_{x}\bigr{)}\;+\;\Upsilon(g)\Bigr{]}.$ (11)
Here, $\Upsilon(g)$ is a regularization term, such as the number of nonzero weights in a linear model or the depth of a decision tree, $\mathcal{L}$ is a loss function, and $\Pi_{x}(\tilde{\bm{x}})$ denotes a proximity measure that quantifies the distance from a perturbation $\tilde{\bm{x}}$ to $\bm{x}$.
II A 2 Limitations
@@
A key component of perturbation-based methods is the perturbation baseline, which is intended to represent the absence of a feature. However, defining an appropriate perturbation baseline is non-trivial. Many RS studies simply employ a scalar value $\alpha$, often set to zero, such as in the quantitative evaluation of Kakageorgiou et al. [26]. However, perturbations with a scalar perturbation baseline can yield Out-of-Distribution (OoD)samples [40]. This is problematic because DL models are prone to unreliable predictions when the input does not reside in their learned distribution [41]. An example case is shown in Fig. 2(a), which depicts an RGB image from the DeepGlobe Multi-Label (DG-ML)dataset. A rectangular patch is perturbed with the Occlusion method, using a black perturbation baseline ($\alpha=0$). Despite occluding a part of the class, the prediction of the model for the class agricultural land increased from $f_{c}(\bm{x})=0.90$ on the original input to $f_{c}(\tilde{\bm{x}})=0.99$ after the occlusion. Even when most of the pixels from the original image are perturbed, as shown in Fig. 2(b), the model predicts with high certainty $f_{c}(\tilde{\bm{x}})=0.38$, highlighting the unpredictable behavior introduced by a poorly chosen perturbation baseline.
@@
This challenge is amplified in multispectral RS images, where different spectral bands (e.g., near-infrared or thermal) capture distinct information. Common baselines, such as all-zero vectors, can be unsuitable because certain bands rarely contain zero values in operational conditions. For example, in thermal imagery, a zero-valued perturbation baseline is not meaningful, as it does not reflect any typical temperature reading. Similarly, for a near-infrared band, some level of infrared radiation is almost always present, making an all-zero replacement an inadequate representation of the absence of a feature. Consequently, selecting a meaningful perturbation baseline requires more sophisticated approaches that account for the different spectral bands.
@@
Figure 2: Visualizations of perturbed images using $\alpha=0$ from the DG-ML dataset for the target class $c$ (agricultural land) and its prediction certainty for the original image: $f_{c}(\bm{{x}})=0.9$. a) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.99$: b) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.38$.
Furthermore, the regions within each RS image are associated with different LULC classes, and unlike in natural images, background content does not exist. Let us consider the case where we have a LULC class $c$ present in the image and let $A_{c}(\bm{x})$ represent the total number of pixels in $\bm{x}$ corresponding to class $c$ (which can be interpreted as the discrete spatial extent of class $c$ within the image). If the area covered by the perturbation mask $A(\Omega)$ of the Occlusion method is smaller than $A_{c}(\bm{x})$, the perturbation may fail to obscure a sufficient number of relevant pixels to significantly decrease the prediction certainty. This is depicted in Fig. 2(a), where we see a rectangular region that is occluded, which does not reduce $f_{c}(\bm{\tilde{x}})$. This limitation is less pronounced in natural images, where discriminative features are often unique (e.g., silhouette, object parts). In such cases, occluding typically leads to a noticeable reduction in prediction certainty even if $A_{c}(\bm{x})>A(\Omega)$.
@@
A similar problem occurs for the LIME method, where $\mathcal{P}_{\mathbb{X}}$ partitions $\bm{x}$ in different superpixels and partially occludes them. Here, the prediction remains unchanged, unless most of the superpixels are replaced with the perturbation baseline (see Fig 2(b)), resulting in a linear model trained primarily on positive class predictions. This imbalance causes the estimated weights of the linear model to be unreliable, leading to an inaccurate interpretation of feature importance. Since no single superpixel significantly alters the decision of the model, LIME assigns almost equal importance to all superpixels. However, this relevance is lower than it should be, as the method underestimates the contribution of discriminative features due to the regularization term, potentially leading to wrong interpretations of the reasoning of the model. Additionally, the more pixels are replaced by the perturbation baseline, the higher the proximity measure $\Pi_{x}$ due to the large distance from the original image. As a result, individual pixels are incorrectly attributed as less relevant.
II B Gradient-propagation-based methods
II B 1 Background
Gradient-propagation-based methods utilize gradient information to generate visual explanations for deep neural network predictions. The gradient-propagation-based explanation method we examine is GradCAM [13], which incorporates gradient information flowing into any convolutional layer, typically the last one due to its high-level semantics and spatial information retention. The attribution map for a specific class $c$ is computed by first capturing the gradients of the class score $Y^{c}$ with respect to the activations $A^{p}$ of convolutional layer $l$. The importance weights $\omega^{c}_{p}$ for each feature map are determined by globally averaging the gradients over all spatial locations $(j,k)$:
$\omega^{c}_{p}=\frac{1}{Z}\sum_{j,k}\frac{\partial Y^{c}}{\partial A^{p}_{j,k}}.$ (12)
These weights represent a partial linearization of the network and denote the importance of each feature map $p$ for the target class $c$. A weighted combination of these maps followed by a ReLU ensures that only positive influences on the class prediction are visualized:

### v3

@@ I Introduction
The development of accurate methods for remote sensing (RS) image scene classification is one of the most important research topics in RS. In recent years, deep learning (DL) has emerged as the dominant approach for scene classification problems due to the high capability of DL models to extract and exploit the complex spatial and spectral content of RS images. RS image scenes can contain multiple land-use land-cover (LULC) classes and thus can be simultaneously associated with different class labels (i.e., multi-labels). Accordingly, the existing models are in general capable of assigning multiple LULC class labels to each RS image scene in an archive [1, 2]. Despite their success, DL models often remain enigmatic black boxes, with almost no human understanding of their internal processes [3]. This lack of transparency raises concerns as decision-makers increasingly rely on these black-box systems [4]. New regulations for critical applications, such as the General Data Protection Regulation [5] and the Artificial Intelligence Act [6], require interpretability to ensure transparent decisions. Beyond regulatory concerns, effective interpretability can also improve model performance by addressing issues such as limited or biased training data, outliers, vulnerability to adversarial attacks, and overfitting [4].
@@
To this end, explainable artificial intelligence (xAI) has emerged as a research area focused on revealing the factors behind the predictions of the DL model. Although xAI methods are widely used in the CV community [7], they have only recently gained attention in RS. For a comprehensive overview of xAI in RS, we refer the reader to [8]. The most prominent type of explanation methods for RS scene classification problems are feature attribution methods [8]. These methods assign importance scores to input features based on their contribution to the prediction of a model[9, 10, 11, 12, 13]. While some feature attribution methods have been adapted to RS scene classification[14, 15], most studies use feature attribution methods that were originally designed for image classification tasks on natural images [16, 17, 18].
@@
This raises the question of whether they are applicable to RS image scenes, as they exhibit different characteristics compared to natural images and can even be considered a distinct modality [19]. To answer this question, it is essential to assess the quality of the generated explanations. Existing evaluation methods can be categorized into two groups: 1) qualitative assessments, which include domain expert evaluations or user studies; and 2) quantitative assessments, which employ explanation metrics [20]. Recently, Höhl et al. [8] reviewed approximately 200 RS publications, finding that the vast majority evaluated qualitative, while only 16 studies included quantitative evaluation using explanation metrics. This reliance on qualitative assessment raises concerns, as studies in the CV community have shown that qualitative assessment is insufficient for the robust verification of explanation methods [21, 22].
@@
One of the key reasons for the limited use of quantitative evaluations in RS is the absence of clear guidelines to assess the quality of the explanation in this domain. The need for such guidelines is underscored by the inherent difficulty of quantitative evaluation, since the explanation metrics depend heavily on their parameterization [23, 22]. Moreover, the lack of ground truth for the explanations renders them unverifiable [24], requiring the use of a wide range of metrics for a more comprehensive evaluation [25]. In many RS studies, only a small part of this range is analyzed, often focusing on single aspects, such as localization capability [17], while only a few studies have evaluated a broader variety of metrics [26, 27]. In particular, none of these studies have addressed the fact that most explanation metrics have been originally designed for natural images and can be limited when applied to RS scene classification.
@@
To address these issues, we present the first unified evaluation of the explanation methods and metrics for RS image scene classification. We assess five widely used feature attribution methods: Occlusion [9], LIME [10], LRP [11], DeepLIFT [12] and GradCAM [13]; and evaluate ten explanation metrics spanning five categories: faithfulness (FE [28], IROF [29]), robustness (AS [30], LLE [28]), localization (RRA [31], TKI [32]), complexity (SP [33], CO [34]) and randomization-based (MPRT [35], RL [36]). These methods and metrics are selected for their popularity in RS and to cover a broad spectrum of evaluation criteria. We perform a methodological analysis to examine whether the existing explanation methods and metrics are suitable for the classification of RS images. Then, we perform an extensive experimental evaluation using MetaQuantus [24] to assess the reliability and effectiveness of the explanation metrics under minor perturbations and disruptive changes, supporting our methodological findings with experimental validation.
@@
  • Based on our methodological and experimental findings, we develop practical guidelines to assist RS practitioners in selecting appropriate explanation methods and evaluation metrics for RS image scene classification and recommend suitable explanation methods and metrics. Our pipeline can be reproduced and adapted for different datasets or classification tasks in RS, providing a valuable resource for the RS community.
The remainder of this paper is organized as follows. Section II and Section III provide a methodological evaluation of the effectiveness of explanation methods and explanation metrics for RS image scene classification, respectively. In Section IV we describe the datasets and the experimental setup. The results of our experiments are presented in Section V. Finally, Section VI draws the conclusion of the work.
Figure 1: The evaluation protocol for selecting appropriate explanation methods and metrics for RS image scene classification. It involves generating explanations using various feature attribution methods, estimating their quality using selected metrics, assessing the reliability of the metric with MetaQuantus [24], and using the most reliable metrics to select the most suitable explanation methods.
@@ II Explanation methods for RS image scene classification
In this section, we provide a methodological analysis of the considered explanation methods with respect to the specific challenges of RS images. Most xAI research has focused on natural images, which differ significantly from RS images [19]. In this paper, we examine commonly applied explanation methods in RS scene classification, focusing on methods that generate feature attribution maps, since these methods are most widely used in RS [8]. However, we note that alternative paradigms such as Multiple Instance Learning [37, 38, 39] or counterfactual explanations [40] have also been utilized in RS. Feature attribution methods can be grouped into perturbation-based and backpropagation-based methods [41]. Perturbation methods modify inputs to measure output changes, while backpropagation methods propagate values from outputs to inputs, typically gradients or relevancy scores. We consider three categories: i) perturbation-based, ii) relevancy-propagation-based, and iii) gradient-propagation-based methods.
Our analysis covers five representative methods: 1) two perturbation-based methods (Occlusion [9], LIME [10]); 2) two relevancy-propagation-based methods (LRP [11], DeepLIFT; [12]) and 3) one gradient-propagation-based method (GradCAM [13]). For each category, we first provide a general description, followed by a definition of the explanation methods and a discussion of the limitations considering the differences between RS and natural images.
Let us consider a RS scene classification problem, where a DL model $f$ is trained on an archive $\bm{X}_{\mathrm{tr}}$ of $N$ pairs $(\bm{x}_{i},\bm{y}_{i})_{i=1}^{N}$ for multi-label classification (the single-label case is omitted for simplicity). Each image $\bm{x}\in\mathbb{R}^{H\times W\times C}$ represents a scene captured by an airborne or satellite sensor, with $C$ spectral bands and $H,W$ spatial dimensions. The model $f$ maps an input image to predicted labels $f(\bm{x})=\hat{\bm{y}}$, where $\hat{\bm{y}}$ represents the predicted class probabilities or binary assignments. To interpret the reasoning behind the prediction of a model $f$ for an input $\bm{x}$, post-hoc explanation methods, such as feature attribution methods $\Phi$, can be applied:
@@
II.A Perturbation-based methods
II.A.1 Background
Perturbation-based methods explain predictions by measuring output changes between original and perturbed inputs. Let $\mathbb{X}\subseteq\mathbb{R}^{H\times W\times C}$ denote the input space and $\Theta$ a set of perturbation parameters. A perturbation function $\mathcal{P}_{\mathbb{X}}:\mathbb{X}\times\Theta\;\to\;\mathbb{X}$ is defined as:
$\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)\;=\;\tilde{\bm{x}}$ (2)
with $\bm{x}\in\mathbb{X}$ and $\theta\in\Theta$. In other words, $\mathcal{P}_{\mathbb{X}}(\bm{x},\theta)$ modifies $\bm{x}$ according to the perturbation parameter $\theta$, while still preserving the overall structure such that $\tilde{\bm{x}}$ is a valid point in the same space. Common perturbation functions are blurring [42], introducing noise [30], or replacing parts of the image [43]. We consider Occlusion [9] and LIME [10] as perturbation-based methods.
In the Occlusion [9] method, each $\theta\in\Theta$ specifies a region (mask) $\Omega\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ and a baseline $\alpha\in\mathbb{R}$. The perturbation function is:
@@ II.A.1 Background
The LIME method $\Phi_{\mathrm{LIME}}$ explains a black-box model $f$ by locally approximating it with an interpretable model trained on binary-encoded perturbations of the input. Let $S=\{S_{1},S_{2},\dots,S_{M}\}$ be a segmentation of $\bm{x}$ into $M$ superpixels. We define $\bm{z}\in\{0,1\}^{M}$ as a binary vector indicating the presence (1) or absence (0) of superpixels. The perturbation function is:
@@
where $\alpha\in\mathbb{R}^{C}$ is a constant baseline. Varying $\bm{z}_{m}$ generates perturbed samples around $\bm{x}$. Then, we fit a simple model $g$ that approximates $f$ near $\bm{x}$:
@@
$\Phi_{\mathrm{LIME}}(\bm{x})\;=\;\operatorname{argmin}_{g\in\mathcal{G}}\Bigl[\mathcal{L}\bigl(f,g,\Pi_{x}\bigr)\;+\;\Upsilon(g)\Bigr],$ (8)
[…]
II.A.2 Limitations
@@
A key component of perturbation-based methods is the perturbation baseline, which represents the absence of the occluded input feature. In practice, many RS studies use a scalar value $\alpha$, often zero, as in Kakageorgiou et al. [26]. However, scalar baselines can produce Out-of-Distribution (OoD) samples [44], and DL models yield unreliable predictions on inputs outside their training distribution [45]. Fig. 2(a) illustrates this limitation with an RGB image from the DeepGlobe Multi-Label (DG-ML) dataset [46]. A rectangular patch is perturbed with the Occlusion method, using a black perturbation baseline ($\alpha=0$). Despite occluding a part of the class, the prediction of the model for the class agricultural land increased from $f_{c}(\bm{x})=0.90$ on the original input to $f_{c}(\tilde{\bm{x}})=0.99$ after the occlusion. Even with most pixels perturbed (Fig. 2(b)), the model outputs $f_{c}(\tilde{\bm{x}})=0.38$, highlighting the unpredictable behavior introduced by a poorly chosen perturbation baseline.
@@
This challenge is amplified in multispectral RS images, where spectral bands capture distinct information. Common baselines, such as all-zero vectors, are unsuitable because some bands rarely take zero values in operational conditions. For example, in thermal images, a zero baseline has no physical meaning, and in near-infrared bands, some radiation is always present. Meaningful baselines are required to account for spectral characteristics.
@@
Figure 2: Visualizations of perturbed images using $\alpha=0$ from the DG-ML [46] dataset for the target class $c$ (agricultural land) and its prediction certainty for the original image: $f_{c}(\bm{{x}})=0.9$. a) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.99$: b) Perturbed sample, where $f_{c}(\bm{\tilde{x}})=0.38$.
[…]
@@
A similar problem occurs for the LIME method, where $\mathcal{P}_{\mathbb{X}}$ partitions $\bm{x}$ in different superpixels and partially occludes them. Unless most superpixels are replaced (Fig. 2(b)), the prediction remains unchanged, resulting in a linear model trained primarily on positive class predictions. This imbalance causes the estimated weights of the linear model to be unreliable, leading to an inaccurate interpretation of feature importance. Since no single superpixel strongly influences the model prediction, LIME assigns nearly equal importance across superpixels, underestimating discriminative features due to regularization. Additionally, the more pixels are replaced by the perturbation baseline, the higher the proximity measure $\Pi_{x}$ due to the large distance from the original image. As a result, individual pixels are incorrectly attributed as less relevant.
II.B Gradient-propagation-based methods
II.B.1 Background
Gradient-propagation-based methods use gradients to explain model predictions. We consider GradCAM [13], which utilizes gradients flowing into a convolutional layer, typically the last one, due to its semantic and spatial content. For class $c$, the gradients of the class score $Y^{c}$ with respect to activations $A^{p}$ of layer $l$ yield importance weights $\omega^{c}_{p}$:
$\omega^{c}_{p}=\frac{1}{Z}\sum_{j,k}\frac{\partial Y^{c}}{\partial A^{p}_{j,k}},$ (9)
where global averaging collapses spatial locations $(j,k)$. These weights approximate the contribution of each feature map $p$ to class $c$. The attribution map is the weighted feature maps followed by ReLU, ensuring only positive contributions are visualized: